OpenStudy (anonymous):

Which of the following best describes the graph of f(x) = x2 - 7x + 10? Minimum at (1.5, -12.25) with intercepts at (5, 0) and (-2, 0) Minimum at (-1.5, -12.25) with intercepts at (-5, 0) and (2, 0) Minimum at (-3.5, -2.25) with intercepts at (-5, 0) and (-2, 0) Minimum at (3.5, -2.25) with intercepts at (5, 0) and (2, 0)

OpenStudy (whpalmer4):

I just answered this. Why did you close and repost?

OpenStudy (jdoe0001):

find the vertex, as shown by @whpalmer4 and then find the x-intercepts by setting y = 0

OpenStudy (anonymous):

so it would be d >?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

what did you get for the vertex?

OpenStudy (jdoe0001):

\(\bf \textit{vertex of a parabola}\\ \quad \\ y = {\color{red}{ a}}x^2+{\color{blue}{ b}}x+{\color{green}{ c}}\qquad \qquad \quad \left(-\cfrac{{\color{blue}{ b}}}{2{\color{red}{ a}}}\quad ,\quad {\color{green}{ c}}-\cfrac{{\color{blue}{ b}}^2}{4{\color{red}{ a}}}\right)\)

OpenStudy (jdoe0001):

to find the x-intercepts -> \(\bf y= x^2 - 7x + 10\implies 0=x^2 - 7x + 10\) solve for "x"

OpenStudy (whpalmer4):

yes, the vertex is at \[x = -\frac{-7}{2*1} = 3.5\] so D is the correct answer. Solving for the x-intercepts, \[x^2-7x+10 = 0\]Factor that as \[(x-5)(x-2) = 0\]\[x-5=0\]\[x=5\]\[x-2=0\]\[x=2\]Intercepts are at \((5,0),(2,0)\)